If 1.5 A flows through R2...

A. what is the potential difference supplied by the ideal battery?

B. what is the voltage drop over R2?

C. what is the current through R3?

If 1.5 A flows through R2... A. what is the potent
Answer #1

A)
Let Inet is the current thrugh battery

current through R2, I2 = Inet*R3/(R2+R3)


1.5 = Inet*20/(20+20)

==> Inet = 1.5*40/20

= 3 A

net resistance of the ckt

Rnet = R1 + R2*R3/(R2+R3) + R4

= 10 + 20*20/(20+20) + 30

= 50 ohms

Apply, V = Inet*Rnet

= 3*50

= 150 volts <<<<<<<<<<<-----------Answer

B) voltage drop over R2, V2 = I2*R2

= 1.5*20

= 30 volts <<<<<<<<<<<-----------Answer


C) current through R3, I3 = Inet*R2/(R2+R3)

= 3*20/(20+20)

= 1.5 A <<<<<<<<<<<-----------Answer